The number of quadratic equations ax2+bx+c=0 such that ax2+bx+c+k=0 has two distinct positive roots for each k,where k is real

Grade:Upto college level

1 Answers

askIITiansexpert nagesh
14 years ago

Dear sudarshan gupta, 

Given, ax2+bx+c+k=0, has two distinct positive roots for each k, where k is real

let, p,q be those  roots

then p and q are distinct and positive, hence product should be always positive for all real k

but pq=(c+k)/a, which can take both positive and negative values as k changes.

So there will be no quadratic equation ax2+bx+c = 0, such that ax2+bx+c+k=0 has 2 distinct positive roots for each real k.

 

 

Please feel free to post as many doubts on our discussion forum as you can. If you find any question

Difficult to understand - post it here and we will get you the answer and detailed solution very quickly. We

are all IITians and here to help you in your IIT JEE preparation. All the best sudarshan gupta

 

Regards,

Askiitians Experts

nagesh